# Representation Learning for Distributional Perturbation Extrapolation

### v1

@@ 8 Discussion
We consider the requirement for the decoder to generalize to new inputs the biggest open problem for distributional perturbation extrapolation. While this issue is absent from our theory, where $\mathbb{P}_{\mathbf{Z}}$ is Gaussian and thus has full support, it can pose major challenges in practice when learning from finite data. As discussed at the end of § 7, the issue of decoder extrapolation is orthogonal to learning the correct representation and perturbation model and thus appears fundamental. If the distribution of perturbed test latents lies fully or partially outside the training support, quantifying the uncertainty associated with different parts of the predicted distribution in observation space would be desirable, e.g., by measuring the overlap in supports or relying on methods for outlier/anomaly detection.
Future Work.
Besides addressing, or at least quantifying, (the lack of) decoder extrapolation, future work should evaluate the proposed approach on more complex, noisy real-world data; extend our theoretical results to partially identifiable settings; and pursue extensions that incorporate covariates and nonlinear dose-response functions.
@@ Appendix A Proofs
A.1 Proof of Thm. 4.1
@@
$\mathbf{Z}_{e}:=\mathbf{Z}+\mathbf{W}\mathbf{l}_{e}$ (A.1)
@@
$\widetilde{\mathbf{Z}}_{e}:=\widetilde{\mathbf{Z}}+\widetilde{\mathbf{W}}\mathbf{l}_{e},$ (A.2)
@@
$p_{e}(\mathbf{z})$ $=p\left(\mathbf{z}-\mathbf{W}\mathbf{l}_{e}\right),$ (A.6)
$\widetilde{p}_{e}(\mathbf{z})$ $=\widetilde{p}\left(\mathbf{z}-\widetilde{\mathbf{W}}\mathbf{l}_{e}\right).$ (A.7)
@@
$\log p(\mathbf{z}-\mathbf{W}\mathbf{l}_{e})-\log p(\mathbf{z}-\mathbf{W}\mathbf{l}_{0})=\log\widetilde{p}\left(\mathbf{h}(\mathbf{z})-\widetilde{\mathbf{W}}\mathbf{l}_{e}\right)-\log\widetilde{p}\left(\mathbf{h}(\mathbf{z})-\widetilde{\mathbf{W}}\mathbf{l}_{0}\right).$ (A.8)
@@
$(\mathbf{l}_{e}-\mathbf{l}_{0})^{\top}\mathbf{W}^{\top}\bm{\Sigma}^{-1}(\mathbf{z}-\bm{\mu})$ $-\frac{1}{2}(\mathbf{l}_{e}-\mathbf{l}_{0})^{\top}\mathbf{W}^{\top}\bm{\Sigma}^{-1}\mathbf{W}(\mathbf{l}_{e}+\mathbf{l}_{0})$ (A.9)
=\,= $(\mathbf{l}_{e}-\mathbf{l}_{0})^{\top}\widetilde{\mathbf{W}}^{\top}\widetilde{\bm{\Sigma}}^{-1}\left(\mathbf{h}(\mathbf{z})-\widetilde{\bm{\mu}}\right)$ $-\frac{1}{2}(\mathbf{l}_{e}-\mathbf{l}_{0})^{\top}\widetilde{\mathbf{W}}^{\top}\widetilde{\bm{\Sigma}}^{-1}\widetilde{\mathbf{W}}(\mathbf{l}_{e}+\mathbf{l}_{0})$ (A.10)
@@
$(\mathbf{l}_{e}-\mathbf{l}_{0})^{\top}\mathbf{W}^{\top}\bm{\Sigma}^{-1}=(\mathbf{l}_{e}-\mathbf{l}_{0})^{\top}\widetilde{\mathbf{W}}^{\top}\widetilde{\bm{\Sigma}}^{-1}\mathbf{J}_{\mathbf{h}}(\mathbf{z}).$ (A.11)
Let $\mathbf{L}\in\mathbb{R}^{K\times M}$ be the matrix with columns $\mathbf{l}_{e}-\mathbf{l}_{0}$ for $e\in[M]$. Then, for all $\mathbf{z}$:
$\mathbf{L}^{\top}\mathbf{W}^{\top}\bm{\Sigma}^{-1}=\mathbf{L}^{\top}\widetilde{\mathbf{W}}^{\top}\widetilde{\bm{\Sigma}}^{-1}\mathbf{J}_{\mathbf{h}}(\mathbf{z}).$ (A.12)
@@
$\bm{0}=\mathbf{L}^{\top}\widetilde{\mathbf{W}}^{\top}\widetilde{\bm{\Sigma}}^{-1}\mathbf{H}_{\mathbf{h}}(\mathbf{z}),$ (A.13)
@@
$\bm{0}=\mathbf{L}^{\top}\widetilde{\mathbf{W}}^{\top}\widetilde{\bm{\Sigma}}^{-1}\frac{\partial^{2}}{\partial z_{i}\partial z_{j}}\mathbf{h}(\mathbf{z}).$ (A.14)
By assumption (iii), the matrix $\mathbf{L}^{\top}\widetilde{\mathbf{W}}^{\top}$ has full column rank and thus a left inverse, i.e., there exists $\mathbf{V}\in\mathbb{R}^{{d_{Z}}\times M}$ such that $\mathbf{V}\mathbf{L}^{\top}\widetilde{\mathbf{W}}^{\top}=\mathbf{I}_{{d_{Z}}}$. Multiplying (A.14) on the left by $\widetilde{\bm{\Sigma}}\mathbf{V}$ then yields for all $i,j\in[{d_{Z}}]$ and all $\mathbf{z}$:
@@
This implies that $\mathbf{h}$ must be affine, i.e., there exist $\mathbf{A}\in\mathbb{R}^{{d_{Z}}\times{d_{Z}}}$ and $\mathbf{b}\in\mathbb{R}^{d_{Z}}$ such that for all $\mathbf{z}$:
$\mathbf{h}(\mathbf{z})=\mathbf{A}\mathbf{z}+\mathbf{b}.$ (A.16)
Further, since $\mathbf{h}$ is invertible, $\mathbf{A}$ must be invertible.
[…]
$\mathcal{N}\left(\widetilde{\bm{\mu}}+\widetilde{\mathbf{W}}\mathbf{l}_{e},\widetilde{\bm{\Sigma}}\right)=\mathcal{N}\left(\mathbf{A}(\bm{\mu}+\mathbf{W}\mathbf{l}_{e})+\mathbf{b},\mathbf{A}\bm{\Sigma}\mathbf{A}^{\top}\right).$ (A.17)
@@
$\mathbf{A}\bm{\Sigma}\mathbf{A}^{\top}=\widetilde{\bm{\Sigma}}\qquad\implies\qquad\mathbf{A}=\widetilde{\bm{\Sigma}}^{\frac{1}{2}}\bm{\Sigma}^{-\frac{1}{2}},$ (A.18)
where the matrix square roots exist and are unique since $\bm{\Sigma}$ and $\widetilde{\bm{\Sigma}}$ are symmetric and positive definite. By equating the means, we obtain for all $e$:
$\widetilde{\bm{\mu}}+\widetilde{\mathbf{W}}\mathbf{l}_{e}=\mathbf{A}(\bm{\mu}+\mathbf{W}\mathbf{l}_{e})+\mathbf{b}\qquad\iff\qquad(\widetilde{\mathbf{W}}-\mathbf{A}\mathbf{W})\mathbf{l}_{e}=\mathbf{A}\bm{\mu}+\mathbf{b}-\widetilde{\bm{\mu}}.$ (A.19)
By contrasting (A.19) for all $e\in[M]$ with $e=0$ as before, we obtain
$(\widetilde{\mathbf{W}}-\mathbf{A}\mathbf{W})\mathbf{L}=\bm{0}\qquad\iff\qquad\widetilde{\mathbf{W}}\mathbf{L}=\mathbf{A}\mathbf{W}\mathbf{L}.$ (A.20)
Finally, by choosing $e=0$ in (A.19) we obtain the desired expression for $\mathbf{b}$,
$\mathbf{b}=\widetilde{\bm{\mu}}-\mathbf{A}\bm{\mu}+(\widetilde{\mathbf{W}}-\mathbf{A}\mathbf{W})\mathbf{l}_{0}.$ (A.21)
@@
If $\mathrm{rank}(\mathbf{L})=K$, then $\mathbf{L}$ has a right inverse, i.e., there exists $\mathbf{K}\in\mathbb{R}^{M\times K}$ such that $\mathbf{L}\mathbf{K}=\mathbf{I}_{K}$. Right multiplication of (A.20) by $\mathbf{K}$ then yields
$\widetilde{\mathbf{W}}=\mathbf{A}\mathbf{W}$ (A.22)
Finally, substitution of (A.22) into (A.21) yields
$\mathbf{b}=\widetilde{\bm{\mu}}-\mathbf{A}\bm{\mu}.$ (A.23)
A.3 Proof of Thm. 4.6
See 4.6

### v2

@@ 8 Discussion, limitations, and future work
Decoder extrapolation. The need for the decoder to generalize to new inputs poses a challenge for distributional perturbation extrapolation. While this issue is absent from our theory, where $\mathbb{P}_{\mathbf{Z}}$ is Gaussian and thus has full support, it arises in practice when learning from finite data. As discussed in Fig.˜3, this issue is orthogonal to learning the correct representation and perturbation model. However, it should be possible to realize that the distribution of perturbed test latents lies (partially) outside the training support. Thus, quantifying the uncertainty associated with different parts of the predicted distribution in observation space may become possible, e.g., by measuring the overlap in supports or relying on outlier and anomaly detection methods.
Modeling gene expression. Our results from Fig.˜4 that methods become indistinguishable for low signal-to-noise ratio are consistent with similar findings in the context of noisy high-dimensional gene perturbation data (Ahlmann-Eltze et al., 2025; Viñas Torné et al., 2025), highlighting the need for better calibrated and interpretable evaluation metrics (Miller et al., 2025; Mejia et al., 2025). Future work should seek to more explicitly incorporate cell-type variation, the zero-inflated nature of gene expression data, and the limited effectiveness of CRISPR perturbations (Hugi et al., 2025) into PDAE and advance our theoretical understanding of extrapolation in the presence of nonlinear interactions.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Proofs
A.1 Proof of reparametrizability without loss of generality
We first state and prove the following lemma, which shows that the choices of means and covariances of the base distribution in assumption (ii) of Thm.˜4.1 come without loss of generality (w.l.o.g.).
Lemma A.1.
Let $\mathbf{f}$ and $\mathbf{W}$ be such that assumptions (i)–(iii) of Thm.˜4.1 are satisfied, and let $\mathbb{P}=\mathcal{N}(\bm{\mu},\bm{\Sigma})$ with positive definite covariance matrix $\bm{\Sigma}$. Then there exist $\widetilde{\mathbf{f}}$ and $\widetilde{\mathbf{W}}$ such that $\widetilde{\mathbf{f}}$, $\widetilde{\mathbf{W}}$, and $\widetilde{\mathbb{P}}=\mathcal{N}(-\widetilde{\mathbf{W}}\mathbf{a}_{0},\mathbf{I})$ generates the same distributions, i.e.,
$\forall e\in[M]_{0}:\qquad\mathbf{f}(\mathbf{Z}+\mathbf{W}\mathbf{a}_{e})\overset{d}{=}\widetilde{\mathbf{f}}(\widetilde{\mathbf{Z}}+\widetilde{\mathbf{W}}\mathbf{a}_{e}),\qquad\text{where}\quad\mathbf{Z}\sim\mathcal{N}(\bm{\mu},\bm{\Sigma})\quad\text{and}\quad\widetilde{\mathbf{Z}}\sim\mathcal{N}(-\widetilde{\mathbf{W}}\mathbf{a}_{0},\mathbf{I}),$ (A.1)
with $\widetilde{\mathbf{f}}$ and $\widetilde{\mathbf{W}}$ also satisfying assumptions (i) and (iii) of Thm.˜4.1.
Proof.
Define $\widetilde{\mathbf{f}}:\mathbf{z}\mapsto\mathbf{f}(\bm{\mu}+\mathbf{W}\mathbf{a}_{0}+\bm{\Sigma}^{\frac{1}{2}}\mathbf{z})$ and let $\widetilde{\mathbf{W}}:=\bm{\Sigma}^{-\frac{1}{2}}\mathbf{W}$. The equality in distribution in (A.1) then follows directly by substitution from the properties of linear transformations of Gaussians. Since $\mathbf{f}$ is a $C^{2}$ diffeomorphism by assumption, so is $\widetilde{\mathbf{f}}$, since it is the composition of $\mathbf{f}$ with an affine function. Moreover, since $\bm{\Sigma}$ (and thus also any inverse square root $\bm{\Sigma}^{-\frac{1}{2}}$) is invertible, we have $\mathrm{rank}(\widetilde{\mathbf{W}}\mathbf{A})=\mathrm{rank}(\mathbf{W}\mathbf{A})$ and thus asssumption (iii) also holds. ∎
A.2 Proof of Thm.˜4.1
@@ A.1 Proof of reparametrizability without loss of generality
$\mathbf{Z}_{e}:=\mathbf{Z}+\mathbf{W}\mathbf{a}_{e}$ (A.2)
@@
$\widetilde{\mathbf{Z}}_{e}:=\widetilde{\mathbf{Z}}+\widetilde{\mathbf{W}}\mathbf{a}_{e},$ (A.3)
@@
$p_{e}(\mathbf{z})$ $=p\left(\mathbf{z}-\mathbf{W}\mathbf{a}_{e}\right),$ (A.7)
$\widetilde{p}_{e}(\mathbf{z})$ $=\widetilde{p}\left(\mathbf{z}-\widetilde{\mathbf{W}}\mathbf{a}_{e}\right).$ (A.8)
@@
$\log p(\mathbf{z}-\mathbf{W}\mathbf{a}_{e})-\log p(\mathbf{z}-\mathbf{W}\mathbf{a}_{0})=\log\widetilde{p}\left(\mathbf{h}(\mathbf{z})-\widetilde{\mathbf{W}}\mathbf{a}_{e}\right)-\log\widetilde{p}\left(\mathbf{h}(\mathbf{z})-\widetilde{\mathbf{W}}\mathbf{a}_{0}\right).$ (A.9)
@@
$(\mathbf{a}_{e}-\mathbf{a}_{0})^{\top}\mathbf{W}^{\top}(\mathbf{z}+\mathbf{W}\mathbf{a}_{0})$ $-\frac{1}{2}(\mathbf{a}_{e}-\mathbf{a}_{0})^{\top}\mathbf{W}^{\top}\mathbf{W}(\mathbf{a}_{e}+\mathbf{a}_{0})$ (A.10)
$=\,$ $(\mathbf{a}_{e}-\mathbf{a}_{0})^{\top}\widetilde{\mathbf{W}}^{\top}\left(\mathbf{h}(\mathbf{z})+\widetilde{\mathbf{W}}\mathbf{a}_{0}\right)$ $-\frac{1}{2}(\mathbf{a}_{e}-\mathbf{a}_{0})^{\top}\widetilde{\mathbf{W}}^{\top}\widetilde{\mathbf{W}}(\mathbf{a}_{e}+\mathbf{a}_{0})$ (A.11)
@@
$(\mathbf{a}_{e}-\mathbf{a}_{0})^{\top}\mathbf{W}^{\top}=(\mathbf{a}_{e}-\mathbf{a}_{0})^{\top}\widetilde{\mathbf{W}}^{\top}\mathbf{J}_{\mathbf{h}}(\mathbf{z}).$ (A.12)
Let $\mathbf{A}\in\mathbb{R}^{K\times M}$ be the matrix with columns $\mathbf{a}_{e}-\mathbf{a}_{0}$ for $e\in[M]$. Then, for all $\mathbf{z}$:
$\mathbf{A}^{\top}\mathbf{W}^{\top}=\mathbf{A}^{\top}\widetilde{\mathbf{W}}^{\top}\mathbf{J}_{\mathbf{h}}(\mathbf{z}).$ (A.13)
@@
$\bm{0}=\mathbf{A}^{\top}\widetilde{\mathbf{W}}^{\top}\mathbf{H}_{\mathbf{h}}(\mathbf{z}),$ (A.14)
@@
$\bm{0}=\mathbf{A}^{\top}\widetilde{\mathbf{W}}^{\top}\frac{\partial^{2}}{\partial z_{i}\partial z_{j}}\mathbf{h}(\mathbf{z}).$ (A.15)
By assumption (iii), the matrix $\mathbf{A}^{\top}\widetilde{\mathbf{W}}^{\top}$ has full column rank and thus a left inverse, i.e., there exists $\mathbf{V}\in\mathbb{R}^{{d_{Z}}\times M}$ such that $\mathbf{V}\mathbf{A}^{\top}\widetilde{\mathbf{W}}^{\top}=\mathbf{I}_{{d_{Z}}}$. Multiplying (A.15) on the left by $\mathbf{V}$ then yields for all $i,j\in[{d_{Z}}]$ and all $\mathbf{z}$:
@@
This implies that $\mathbf{h}$ must be affine, i.e., there exist $\mathbf{M}\in\mathbb{R}^{{d_{Z}}\times{d_{Z}}}$ and $\mathbf{b}\in\mathbb{R}^{d_{Z}}$ such that for all $\mathbf{z}$:
$\mathbf{h}(\mathbf{z})=\mathbf{M}\mathbf{z}+\mathbf{b}.$ (A.17)
Further, since $\mathbf{h}$ is invertible, $\mathbf{M}$ must be invertible.
Recall that $\widetilde{\mathbf{Z}}_{e}\overset{d}{=}\mathbf{h}(\mathbf{Z}_{e})=\mathbf{M}\mathbf{Z}_{e}+\mathbf{b}$. It follows from (A.2), (A.3) and assumption (ii) that for all $e\in[M]_{0}$:
$\mathcal{N}\left(\widetilde{\mathbf{W}}(\mathbf{a}_{e}-\mathbf{a}_{0}),\mathbf{I}\right)=\mathcal{N}\left(\mathbf{M}\mathbf{W}(\mathbf{a}_{e}-\mathbf{a}_{0})+\mathbf{b},\mathbf{M}\mathbf{M}^{\top}\right).$ (A.18)
@@
$\mathbf{M}\mathbf{M}^{\top}=\mathbf{I},$ (A.19)
i.e., $\mathbf{M}=\mathbf{O}$ for some orthogonal matrix $\mathbf{O}\in O({d_{Z}})$.
Similarly, by equating the means, we obtain for all $e$:
$\widetilde{\mathbf{W}}(\mathbf{a}_{e}-\mathbf{a}_{0})=\mathbf{O}\mathbf{W}(\mathbf{a}_{e}-\mathbf{a}_{0})+\mathbf{b}$ (A.20)
For $e=0$, (A.20) yields
$\mathbf{b}=\bm{0}.$ (A.21)
Stacking (A.20) for all other $e\in[M]$, we obtain
$\widetilde{\mathbf{W}}\mathbf{A}=\mathbf{O}\mathbf{W}\mathbf{A}.$ (A.22)
@@
If $\mathrm{rank}(\mathbf{A})=K$, then $\mathbf{A}$ has a right inverse, i.e., there exists $\mathbf{K}\in\mathbb{R}^{M\times K}$ such that $\mathbf{A}\mathbf{K}=\mathbf{I}_{K}$. Right multiplication of (A.22) by $\mathbf{K}$ then yields
$\widetilde{\mathbf{W}}=\mathbf{O}\mathbf{W}.$ (A.23)
A.4 Proof of Thm.˜4.7
See 4.7
@@
$\mathbf{h}(\mathbf{Z}+\mathbf{W}\mathbf{a}_{\mathrm{test}})\overset{d}{=}\widetilde{\mathbf{Z}}+\widetilde{\mathbf{W}}\mathbf{a}_{\mathrm{test}}.$ (A.24)
By Thm.˜4.1, we have $\mathbf{h}(\mathbf{z})=\mathbf{O}\mathbf{z}$ for some orthogonal matrix $\mathbf{O}\in O(d_{Z})$.
Together with $\mathbf{Z}\sim\mathcal{N}(-\mathbf{W}\mathbf{a}_{0},\mathbf{I})$, this lets us compute the distribution of the LHS of (A.24) as:
$\mathbf{h}(\mathbf{Z}+\mathbf{W}\mathbf{a}_{\mathrm{test}})=\mathbf{O}(\mathbf{Z}+\mathbf{W}\mathbf{a}_{\mathrm{test}})\sim\mathcal{N}\left(-\mathbf{O}\mathbf{W}\mathbf{a}_{0}+\mathbf{O}\mathbf{W}\mathbf{a}_{\mathrm{test}},\mathbf{I}\right).$ (A.25)
Similarly, since $\widetilde{\mathbf{Z}}\sim\mathcal{N}\left(-\widetilde{\mathbf{W}}\mathbf{a}_{0},\mathbf{I}\right)$, the distribution of the RHS of (A.24) is given by
$\widetilde{\mathbf{Z}}+\widetilde{\mathbf{W}}\mathbf{a}_{\mathrm{test}}\sim\mathcal{N}\left(-\widetilde{\mathbf{W}}\mathbf{a}_{0}+\widetilde{\mathbf{W}}\mathbf{a}_{\mathrm{test}},\mathbf{I}\right).$ (A.26)
$\mathbf{O}\mathbf{W}\left(\mathbf{a}_{\mathrm{test}}-\mathbf{a}_{0}\right)=\widetilde{\mathbf{W}}\left(\mathbf{a}_{\mathrm{test}}-\mathbf{a}_{0}\right)$ (A.27)
Next, by (4.5), i.e., the assumption that $(\mathbf{a}_{\mathrm{test}}-\mathbf{a}_{0})$ lies in the span of $\{\mathbf{a}_{e}-\mathbf{a}_{0}\}_{e\in[M]}$, there exists $\bm{\alpha}\in\mathbb{R}^{M}$ such that
$\mathbf{a}_{\mathrm{test}}-\mathbf{a}_{0}=\sum_{e\in[M]}\alpha_{e}(\mathbf{a}_{e}-\mathbf{a}_{0})=\mathbf{A}\bm{\alpha}$ (A.28)
where, as before, $\mathbf{A}\in\mathbb{R}^{K\times M}$ is the matrix with columns $(\mathbf{a}_{e}-\mathbf{a}_{0})$. Substituting (A.28) into (A.27) yields
$\mathbf{O}\mathbf{W}\mathbf{A}\bm{\alpha}=\widetilde{\mathbf{W}}\mathbf{A}\bm{\alpha}$ (A.29)
Finally, it follows from Thm.˜4.1 that
$\mathbf{O}\mathbf{W}\mathbf{A}=\widetilde{\mathbf{W}}\mathbf{A}.$ (A.30)
which upon substitution into (A.29) completes the proof. ∎
A.5 Proof of Prop.˜4.9
See 4.9
Proof.
The distribution of $\mathbf{Z}$ induced by the SEM is most easily understood via the reduced form expression
$\mathbf{Z}=(\mathbf{I}-\mathbf{B}^{\top})^{-1}\bm{\eta}$ (A.31)
where the inverse $(\mathbf{I}-\mathbf{B}^{\top})^{-1}$ exists since, by assumption, $\rho(\mathbf{B})<1$. The observational (unintervened) distribution is thus given by $\mathbb{P}_{\mathbf{Z}}=(\mathbf{I}-\mathbf{B}^{\top})^{-1}_{\#}\mathbb{P}_{\bm{\eta}}$. Now consider the shift interventions from (4.7). Analogous to (A.31), the reduced form of (4.7) is given by
$\mathbf{Z}=(\mathbf{I}-\mathbf{B}^{\top})^{-1}\bm{\eta}+(\mathbf{I}-\mathbf{B}^{\top})^{-1}\mathbf{a}_{e}.$ (A.32)
which is equal in distribution to
$\mathbf{Z}^{\mathrm{pert}}=\mathbf{Z}^{\mathrm{base}}+\mathbf{W}\mathbf{a}_{e}$ (A.33)
with $\mathbf{Z}^{\mathrm{base}}\sim\mathbb{P}_{\mathbf{Z}}=(\mathbf{I}-\mathbf{B}^{\top})^{-1}_{\#}\mathbb{P}_{\bm{\eta}}$ and $\mathbf{W}=(\mathbf{I}-\mathbf{B}^{\top})^{-1}$. ∎
Remark A.2.
The correspondence between mean shift perturbations and shift interventions appears to only hold for linear SCMs (at least in this simple form). Consider, for example, a nonlinear additive noise model,
$\mathbf{Z}:=\mathbf{f}(\mathbf{Z})+\bm{\eta},$ (A.34)
with reduced form given by $\mathbf{Z}=\mathbf{g}(\bm{\eta})$, where $\mathbf{g}$ is the inverse of the mapping $\mathbf{z}\mapsto\mathbf{z}-\mathbf{f}(\mathbf{z})$. For shift interventions in (A.34) to match our perturbation model, we then must have
$\mathbf{Z}_{e}=\mathbf{g}(\bm{\eta}+\mathbf{a}_{e})=\mathbf{g}\left(\mathbf{g}^{-1}(\mathbf{Z})+\mathbf{a}_{e}\right)={\bm{\phi}}\left(\mathbf{Z},\mathbf{a}_{e}\right)$ (A.35)
for suitable $\bm{\phi}$} and $\mathbf{a}_{e}$. Thus, if $\mathbf{f}$ is nonlinear, so is $\mathbf{g}$ and this implies under weak assumptions that $\bm{\phi}$} is nonlinear, too. In this case, shift interventions in a nonlinear SCM do not, in general, amount to mean shift perturbations (i.e., linear $\bm{\phi}$}).
Appendix B PDAE training: Additional losses and complete algorithm
In this appendix, we discuss additional loss compoments (other than the perturbation loss from the main text), which are not required based on our theoretical results, but which may still benefit empirical performance, see Fig.˜5 for an illustration.
Reconstructions. When considering the synthetic observations from (5.3) for $h=e$, i.e., when evaluated on the estimated perturbed latents from (5.1), we refer to the corresponding (random) decoder outputs as reconstructions,
$\widehat{\mathbf{X}}_{e,i}=\widehat{\mathbf{f}}\big(\widehat{\mathbf{z}}^{\mathrm{pert}}_{e,i},\,{\bm{\varepsilon}}\big)\qquad\text{where}\qquad{\bm{\varepsilon}}\sim\mathbb{Q}_{\bm{\varepsilon}}.$ (B.1)
Conditional distribution over reconstructions. For any given observation $\mathbf{x}$, we can sample reconstructions from the stochastic decoder by fixing its first argument to the encoding $\widehat{\mathbf{g}}(\mathbf{x})$ and sampling noise from $\mathbb{Q}_{\bm{\varepsilon}}$ as in (B.1). Together, the encoder, decoder, and noise distribution thus induce a distribution over reconstructions, formally given by the push-forward
$\widehat{\mathbf{f}}\left(\widehat{\mathbf{g}}(\mathbf{x}),\cdot\right)_{\#}\mathbb{Q}_{\bm{\varepsilon}}.$ (B.2)
Since this distribution should ideally match that of $\mathbf{X}|\widehat{\mathbf{g}}(\mathbf{X})=\widehat{\mathbf{g}}(\mathbf{x})$ (i.e., the true distribution of observations that share the same encoding), [^3] For example, if $\widehat{\mathbf{g}}$ is injective, the true reconstruction distribution is a point mass on $\mathbf{x}$, whereas if $\widehat{\mathbf{g}}$ is constant, it is the full (unconditional) distribution of $\mathbf{X}$. In this sense, the distribution of reconstructions captures the information about $\mathbf{x}$ that is left unexplained given only its encoding (Shen & Meinshausen, 2024a). we denote it by $\widehat{\mathbb{P}}_{\mathbf{X}|\widehat{\mathbf{g}}(\mathbf{X})=\widehat{\mathbf{g}}(\mathbf{x})}$. Note that the first argument of $\widehat{\mathbf{f}}$ in (B.2) is considered fixed, and all stochasticity stems from the noise $\bm{\varepsilon}$}. This contrasts with the simulated perturbed distributions $\mathbb{P}_{e\to h}$ in (5.4), where both $\mathbf{X}_{e}$ and $\bm{\varepsilon}$} are random.
Conditional reconstruction loss. The conditional reconstruction loss $\mathcal{L}_{\text{cond-rec}}$ is given by the sum of domain-specific DPA losses (Shen & Meinshausen, 2024a),
$\mathcal{L}_{\text{cond-rec}}\left(\widehat{\mathbf{g}},\widehat{\mathbf{f}};\left\{\mathbb{P}_{e}\right\}_{e\in[M]_{0}}\right)=\sum_{e\in[M]_{0}}-\mathbb{E}_{\mathbf{X}_{e}\sim\mathbb{P}_{e}}\left[\operatorname{ES}_{\beta}\left(\widehat{\mathbb{P}}_{\mathbf{X}|\widehat{\mathbf{g}}(\mathbf{X})=\widehat{\mathbf{g}}(\mathbf{X}_{e})},\mathbf{X}_{e}\right)\right],$ (B.3)
where each summand on the RHS of (LABEL:eq:reconstruction_loss_main) involves a conditional energy score (i.e., one whose first argument also depends on $\mathbf{X}_{e}$), see §˜C.3 for further discussion. It should be noted that the pairwise sum in the perturbation loss of (5.5) also includes the case $h=e$. Enforcing $\mathbb{P}_{e}=\widehat{\mathbb{P}}_{e\to e}$ can be viewed as a marginal reconstruction objective and thus as a weaker constraint than that enforced by $\mathcal{L}_{\text{cond-rec}}$.
Figure 5: Extended illustration of the PDAE architecture including the reconstruction loss.
Prior loss. To turn PDAE into a full generative model and remove unnecessary ambiguities from the learned representation, we propose an additional loss which quantifies deviations of the encoded latent basal state distributions from a fixed Gaussian prior. Since, by Lemma˜A.1, the prior mean and covariance are arbitrary, we choose $\mathbb{P}^{\mathrm{base}}=\mathcal{N}(\bm{0},\mathbf{I})$ for convenience. Denote by $\widehat{\mathbb{P}}_{\widehat{\mathbf{g}}(\mathbf{X}_{e})-\mathbf{W}\mathbf{a}_{e}}$ the distribution of
$\widehat{\mathbf{g}}(\mathbf{X}_{e})-\widehat{\mathbf{W}}\mathbf{a}_{e},\qquad\text{where}\qquad\mathbf{X}_{e}\sim\mathbb{P}_{e}.$ (B.4)
We then define the prior loss as the sum over domains of the negative expected energy score between the true fixed prior and the encoded basal state distribution from domain $e$,
$\mathcal{L}_{\text{prior}}\left(\widehat{\mathbf{g}},\widehat{\mathbf{W}};\left\{\left(\mathbb{P}_{e},\mathbf{a}_{e}\right)\right\}_{e=0}^{M}\right):=\sum_{e\in[M]_{0}}-\mathbb{E}_{\mathbf{Z}^{\mathrm{base}}\sim\mathbb{P}^{\mathrm{base}}}\left[\operatorname{ES}_{\beta}\left(\widehat{\mathbb{P}}_{\widehat{\mathbf{g}}(\mathbf{X}_{e})-\widehat{\mathbf{W}}\mathbf{a}_{e}},\mathbf{Z}^{\mathrm{base}}\right)\right].$ (B.5)
In practice, the prior loss can be approximated using the estimated latent basal states,
$\widehat{\mathbf{z}}^{\mathrm{base}}_{e,i}:=\widehat{\mathbf{z}}^{\mathrm{pert}}_{e,i}-\widehat{\mathbf{W}}\mathbf{a}_{e}=\widehat{\mathbf{g}}(\mathbf{x}_{e,i})-\widehat{\mathbf{W}}\mathbf{a}_{e}$ (B.6)
Sparsity. Similar to Bereket & Karaletsos (2023), we also consider a sparsity penalty $\left\lVert\widehat{\mathbf{W}}\right\rVert_{2,1}$ on the inferred $\widehat{\mathbf{W}}$.
Combined loss. Given loss weight hyperparameters $(\lambda_{\textsc{r}},\lambda_{\text{prior}},\lambda_{\textsc{s}})$ and denoting the perturbation loss from (5.5) by $\mathcal{L}_{\mathrm{pert}}\left(\widehat{\mathbf{g}},\widehat{\mathbf{f}},\widehat{\mathbf{W}};\left\{\left(\mathbb{P}_{e},\mathbf{a}_{e}\right)\right\}_{e=0}^{M}\right)$ we can then combine the different losses into one objective as follows:
$\left(\widehat{\mathbf{g}}^{*},\widehat{\mathbf{f}}^{*},\widehat{\mathbf{W}}^{*}\right)\in\operatorname*{arg\,min}_{\widehat{\mathbf{g}},\widehat{\mathbf{f}},\widehat{\mathbf{W}}}$ $\mathcal{L}_{\mathrm{pert}}\left(\widehat{\mathbf{g}},\widehat{\mathbf{f}},\widehat{\mathbf{W}};\left\{\left(\mathbb{P}_{e},\mathbf{a}_{e}\right)\right\}_{e=0}^{M}\right)+\lambda_{\textsc{r}}\mathcal{L}_{\text{cond-rec}}\left(\widehat{\mathbf{g}},\widehat{\mathbf{f}};\left\{\mathbb{P}_{e}\right\}_{e=0}^{M}\right)$ (B.7)
$+\lambda_{\text{prior}}\mathcal{L}_{\text{prior}}\left(\widehat{\mathbf{g}},\widehat{\mathbf{W}};\left\{\left(\mathbb{P}_{e},\mathbf{a}_{e}\right)\right\}_{e=0}^{M}\right)+\lambda_{\textsc{s}}\left\lVert\widehat{\mathbf{W}}\right\rVert_{2,1}.$
To avoid that the encoder captures latent dimensions of maximal variation as in DPA (Shen & Meinshausen, 2024a) rather than focus on the perturbation-relevant latents $\mathbf{Z}$, we do not update the encoder parameters using the reconstruction loss.
Algorithm. The complete algorithm for one training iteration for PDAE is provided in Alg.˜1.
Algorithm 1 PDAE training step